# Utilizing smartphone sensors for accurate solar irradiance measurement and educational purposes

### v2

@@ 6 Conclusion
This study reveals the remarkable potential of smartphone light sensors as effective tools for measuring direct solar irradiance and for introducing students to the fundamental aspects of the solar resource within instructional laboratories. The proposed experiment involves aligning the equipment with the Sun at all times using a home-made collimator tube, thereby allowing to focus on the smartphone’s ability to measure irradiance without considering geometrical and optical aspects that influence other solar irradiance measurements (i.e. global horizontal). In terms of equipment positioning while this study employed a professionally precise solar tracker to gather continuous measurements throughout the day, for instructional laboratory purposes students can manually perform measurements at regular intervals throughout a day. This modification enables hands-on student engagement and participation.
@@
Two calibration methods were presented to convert smartphone illuminance measurements into solar irradiance: comparison with a reference measurement and utilization of clear-sky satellite estimates, which are readily available for any location on the globe (the McClear model). Remarkably, this work demonstrated that calibration against both data sets can be performed without introducing a significant increase in uncertainty, resulting in highly reliable measurement capabilities suitable for instructional laboratories. It is crucial to conduct calibration on a clear-sky day, following the same principles applied in professional calibrations according to current ISO standards for commercial radiometers. The findings of this study pave the way for the development of various low-cost instructional laboratories, both within traditional classroom settings and in outdoor environments. In the process, students must learn to develop important scientific-mathematical skills such as the recording and processing of experimental data, assess their quality with or without data filtering, and assigning uncertainty to simplified models for direct solar irradiance estimation at ground level. Moreover, the calibration process itself allows students to be introduced to simple linear regression, and the uncertainty assignment to its parameters and the final result.
This work shows that smartphones are a high-potential measurement and learning tool, as well as a laboratory in the pocket of every citizen, also on issues related to solar energy. This is a current and relevant topic to be introduced in scientific-technological university careers, given the energy transition that humanity must go through in the coming decades from a high energy dependence on fossil fuels towards renewable energies. The demonstration that smartphones are capable of measuring solar irradiance with low uncertainty and reporting verifiable results with professional measurements and estimates, opens up several didactic possibilities and future studies. To mention a few, it is possible to consider the measurement of global irradiance (direct + diffuse) in a horizontal plane and analyze the directional response of the smartphone sensor, carry out studies to assign a typical measurement uncertainty to the equipment in relation to professional equipment and/or study the behavior of the sensor spectra, using spectral measurements of solar irradiance. By leveraging smartphone technology, students can gain practical insights into solar irradiance measurement, fostering a deeper understanding of this important aspect of renewable energy resources.
Acknowledgements
We thank Sebastian Staacks, developer of PhyPhox, who especially made a more accurate version of the app. The authors have no conflicts to disclose.
Data availability

### v3

@@ 5 Conclusion
This study reveals the remarkable potential of smartphone light sensors as effective tools for measuring direct solar irradiance and for introducing students to the fundamental aspects of the solar resource. To validate the possibility of calibrating smartphones, this experiment was conducted using professional solar measurement and tracking equipment t Once this feasibility was demonstrated, instructors could guide students to prepare their own smartphones similarly to the experiment, using tracing paper and a simple collimator tube, to measure direct illuminance in normal incidence. These measurements, along with direct solar irradiance data measured on-site or from satellite estimates, could be used to accurately calibrate the smartphones, providing an innovative tool to advance understanding of solar resources. Remarkably, this work demonstrated that calibration against both data sets can be performed without introducing a significant increase in uncertainty, resulting in highly reliable measurement capabilities suitable for instructional laboratories. It is important to conduct calibration on a clear-sky day, following the same principles applied in professional calibrations according to current ISO standards for commercial radiometers.
@@
Our findings pave the way for the development of various low-cost instructional laboratories, both within traditional classroom settings and in outdoor environments. In the process, students must learn to develop important scientific-mathematical skills such as the recording and processing of experimental data, assess their quality with or without data filtering, and assigning uncertainty to simplified models for direct solar irradiance estimation at ground level. By leveraging smartphone technology, students can gain practical insights into solar irradiance measurement, fostering a deeper understanding of this important aspect of renewable energy resources.
Conceptualization, J.L.D.L. and R.A.S.; methodology, J.L.D.L. and R.A.S.; software, A.M. ; validation, J.L.D.L. and R.A.S and M.M.; data curation, J.L.D.L. and M.M.; writing—original draft preparation, J.L.D.L.; R.A.S and A.C.M.; writing—review and editing, J.L.D.L.; R.A.S and A.C.M.; visualization, J.L.D.L. and M.M. All authors have read and agreed to the published version of the manuscript.
This research received no external funding.
Not applicable.
Not applicable.
@@
Acknowledgements.
[…]
DOAJ | Directory of open access journals |
TLA | Three letter acronym |
LD | Linear dichroism |